OpenStudy (anonymous):

The length of time t a satellite takes to complete a circular orbit of Earth varies directly as the radius r of the orbit and inversely as the orbital velocity v of the satellite. If t = 1.42 hours when r = 4,050 miles and v = 18,000 miles per hour, find t for r = 4,300 miles and v = 18,500 miles per hour. Round to two decimals.

OpenStudy (ybarrap):

$$ t={\alpha~r\over v} $$ Let's find \(\alpha\) $$ 1.42={\alpha ~4,050\over 18,000}\\ \implies \alpha={1.42\times18,000\over 4,050} $$ For r=4,300 and v=18,500 $$ t={\alpha~4,300 \over 18,500} $$ Does this make sense?

OpenStudy (anonymous):

Yes thank you

OpenStudy (ybarrap):

You're welcome